Service calls arriving at an electric company follow a Poisson distribution with an average arrival rate of 59 per hour. Using the normal approximation to the Poisson, find the probability that the electric company receives at most 49 service calls per hour. Round your answer to four decimal places, if necessary.
Answer #1

Mean = 59

Standard deviation = = 7.6811

Hence,

P(Atmost 49)

= P(X < 49.5) [Continuity Correction]

= P(z < (49.5 - 59)/7.6811)

= P(z < -1.24)

= 0.1081